OpenStudy (anonymous):

1. What is the magnitude of the force on an electron moving with a velocity v = 6.33106 m/s at right angles to a magnetic field B = 2.5310-2 T?

OpenStudy (nurali):

F= e(V x B) e=1.602×10^-19 V=6.33*10^6 m/s B = 2.53*10^-2 T Solve it F=?

OpenStudy (anonymous):

Thank. you. :)

OpenStudy (nurali):

My Pleasure.

OpenStudy (anonymous):

Well, my graphic calculator ran out of power so now I have to wait. :/

OpenStudy (nurali):

Ok.

OpenStudy (anonymous):

Is that Harry Potter in your profile or you? lol

OpenStudy (nurali):

Harry Potter Picture

OpenStudy (anonymous):

I got 2.565=F